# The algebraic and geometric classification of

### v1

@@ 2.2. The geometric classification of compatible commutative associative algebras
Theorem 19.
The variety of complex $2$-dimensional compatible commutative associative algebras has dimension $7$ and it has $2$ irreducible components defined by $\overline{\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})}$ and $\overline{\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})}.$ In particular, there are no rigid algebras in this variety.
@@
Thanks to Theorem 8, we have the algebraic classification of $2$-dimensional compatible commutative associative algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})$ $=$ $7$  and $\dim\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})$ $=$ $6$
$\mathcal{C}_{38}^{*}\not\to\mathcal{C}_{39}^{\alpha,\beta}$ due to the following relation
=$\mathcal{R}=\left\{c_{22}^{1}=c_{12}^{1}=0,\ c_{22}^{2}c_{12}^{\prime 2}+c_{12}^{2}c_{21}^{\prime 1}+c_{11}^{1}c_{22}^{\prime 2}=c_{11}^{\prime 1}c_{22}^{2}+c_{11}^{1}c_{21}^{\prime 1}+c_{12}^{2}c_{22}^{\prime 2}\right\}.$
Thanks to [27], we have $\mathcal{C}_{07}\to\{\mathcal{C}_{03},\mathcal{C}_{05}^{0},\mathcal{C}_{06}^{1}\}.$ All necessary degenerations are given below
$\mathcal{C}_{38}^{0,0,t^{-1}}\xrightarrow{(te_{1},\ te_{2})}\mathcal{C}_{07},$   $\mathcal{C}_{39}^{-it^{-1},it^{-1}}\xrightarrow{(ite_{1}-ite_{2},\ -t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{13}^{1},$  $\mathcal{C}_{38}^{0,0,\alpha}\xrightarrow{(ite_{1}-ite_{2},\ -t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{15}^{\alpha},$   $\mathcal{C}_{38}^{\frac{i}{2t},0,\frac{i}{t}}\xrightarrow{(ite_{1}-ite_{2},\ -t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{16}^{0},$  $\mathcal{C}_{38}^{-\frac{1+\alpha t^{2}}{4t^{2}},-\frac{1+\alpha t^{2}}{4t^{2}},\frac{\alpha t^{2}-3}{4t^{2}}}\xrightarrow{(ite_{1}-ite_{2},\ -t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{18}^{\alpha},$  $\mathcal{C}_{38}^{0,-t,\beta+t}\xrightarrow{(e_{1},\ te_{2})}\mathcal{C}_{24}^{0,\beta,0},$  $\mathcal{C}_{38}^{0,0,\beta}\xrightarrow{(e_{1},\ te_{2})}\mathcal{C}_{25}^{0,\beta,0},$  $\mathcal{C}_{39}^{\alpha,t^{-1}}\xrightarrow{(e_{1},\ te_{2})}\mathcal{C}_{29}^{\alpha},$  $\mathcal{C}_{39}^{-t+\beta,\beta}\xrightarrow{(e_{1}+e_{2},\ te_{2})}\mathcal{C}_{31}^{1,\beta,\beta},$  $\mathcal{C}_{38}^{-\alpha-\beta t+t^{-1},-\beta t,t^{-1}}\xrightarrow{(e_{1}+e_{2},\ te_{2})}\mathcal{C}_{30}^{\alpha,\beta},$  $\mathcal{C}_{38}^{0,\beta,0}\xrightarrow{(e_{1}+e_{2},\ te_{2})}\mathcal{C}_{32}^{1,\beta,\beta},$  $\mathcal{C}_{38}^{-t^{-2},\alpha-(1+\beta t)t^{-2},(1+\beta t)t^{-2}}\xrightarrow{(e_{1}+e_{2},\ te_{2})}\mathcal{C}_{34}^{\alpha,\beta},$  $\mathcal{C}_{38}^{\frac{1+2\alpha t}{2t},\alpha,0}\xrightarrow{(e_{1}+e_{2},\ -te_{1}+te_{2})}\mathcal{C}_{35}^{\alpha}.$
2.3. The geometric classification of compatible associative algebras
@@
Thanks to Theorem 10, we have the algebraic classification of $2$-dimensional compatible associative algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})$ $=$ $7,$ $\dim\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})$ $=$ $6,$ $\dim\mathcal{O}(\mathcal{C}_{28})$ $=$ $\dim\mathcal{O}(\mathcal{C}_{34})$ $=$ $4$
Thanks to Theorem 20, we have that $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{39}^{\alpha,\beta}$ give irreducible components in the variety of $2$-dimensional compatible commutative associative algebras. The rest of the necessary degenerations are given below.
$\mathcal{C}_{28}\xrightarrow{(e_{2},\ te_{1})}\mathcal{C}_{05}^{1},$  $\mathcal{C}_{33}\xrightarrow{(e_{2},\ te_{1})}\mathcal{C}_{06}^{0},$  $\mathcal{C}_{28}\xrightarrow{(e_{1}+\beta e_{2},\ te_{2})}\mathcal{C}_{25}^{1,\beta,\beta},$  $\mathcal{C}_{33}\xrightarrow{(e_{1}+\beta e_{2},\ te_{2})}\mathcal{C}_{32}^{0,\beta,0}.$
2.4. The geometric classification of compatible Novikov algebras
@@
Theorem 21.
The variety of complex $2$-dimensional compatible Novikov algebras has dimension $7$ and it has $6$ irreducible components defined by $\overline{\mathcal{O}(\mathcal{C}_{33})},$  $\overline{\mathcal{O}(\mathcal{C}_{09}^{\alpha,\beta})},$ $\overline{\mathcal{O}(\mathcal{C}_{22}^{\alpha,\beta})},$ $\overline{\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})},$ $\overline{\mathcal{O}(\mathcal{C}_{31}^{\alpha,\beta,\gamma})}$ and $\overline{\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})}.$ In particular, there is one rigid algebra in this variety.
@@
Thanks to Theorem 12, we have the algebraic classification of $2$-dimensional compatible Novikov algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})$ $=$ $7,$
$\dim\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{09}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{22}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{31}^{\alpha,\beta,\gamma})$ $=$ $6,$
$\dim\mathcal{O}(\mathcal{C}_{33})$ $=$ $4$
Thanks to Theorem 21, we have that $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{39}^{\alpha,\beta}$ give irreducible components in the variety of $2$-dimensional compatible commutative associative algebras. Algebras $\mathcal{C}_{09}^{\alpha,\beta},$ $\mathcal{C}_{22}^{\alpha,\beta},$ and $\mathcal{C}_{31}^{\alpha,\beta,\gamma}$ have a one-dimensional subalgebra with zero multiplication (concerning both multiplications), but $\mathcal{C}_{33}$ has not. Hence, $\mathcal{C}_{33}$ is not in the orbit closure of $\mathcal{C}_{09}^{\alpha,\beta},$ $\mathcal{C}_{22}^{\alpha,\beta},$ and $\mathcal{C}_{31}.$ The rest of the necessary degenerations are given below.
$\mathcal{C}_{31}^{t,\frac{t+2}{2t},-\frac{1}{2}}\xrightarrow{(te_{1}+\frac{2t}{3}e_{2},\ \frac{t^{2}}{2}e_{1}+t^{2}e_{2})}\mathcal{C}_{01},$  $\mathcal{C}_{31}^{0,0,t^{-1}}\xrightarrow{(t^{2}e_{1}+e_{2},\ te_{1}-t^{2}e_{2})}\mathcal{C}_{04},$  $\mathcal{C}_{31}^{0,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}-\alpha^{-1}t^{2}e_{2},\ te_{2})}\mathcal{C}_{06}^{\alpha\neq 0},$  $\mathcal{C}_{09}^{\frac{(2\alpha+t)\beta}{2\beta+t},-\frac{\beta t}{2\beta+t}}\xrightarrow{(\frac{2\beta+t}{2\beta}e_{1}-\frac{(2\beta+t)^{2}}{2\beta t}e_{2},\ \frac{(2\beta+t)t}{4\beta^{2}}te_{1})}\mathcal{C}_{10}^{\alpha,\beta\neq 0},$  $\mathcal{C}_{31}^{1,\alpha t^{-1},t^{-1}}\xrightarrow{(te_{1}-t^{2}e_{2},\ -t^{3}e_{2})}\mathcal{C}_{13}^{\alpha},$  $\mathcal{C}_{31}^{1,t^{-1},0}\xrightarrow{(te_{1}+\alpha^{-1}te_{2},\alpha^{-1}t^{2}e_{2})}\mathcal{C}_{14}^{\alpha\neq 0},$  $\mathcal{C}_{31}^{t^{-1}-t,0,-1+\alpha t^{-1}}\xrightarrow{(t^{2}e_{1}+\frac{t^{2}}{1-\alpha t}e_{2},\ te_{1}+\frac{t^{3}}{1-\alpha t}e_{2})}\mathcal{C}_{20}^{\alpha,0},$  $\mathcal{C}_{31}^{t^{-1}-t,0,-\alpha t^{-1}}\xrightarrow{(t^{2}e_{1}-\alpha^{-1}te_{2},\ te_{1}-\alpha^{-1}te_{2})}\mathcal{C}_{21}^{\alpha,0},$  $\mathcal{C}_{31}^{t^{-1}-t,\alpha t^{-1},\beta t^{-1}}\xrightarrow{(t^{2}e_{1}-\beta^{-1}te_{2},\ te_{1}-\beta^{-1}t^{2}e_{2})}\mathcal{C}_{23}^{\alpha,\beta},$  $\mathcal{C}_{31}^{\alpha,\beta,\gamma}\xrightarrow{(e_{1},\ t^{-1}e_{2})}\mathcal{C}_{32}^{\alpha,\beta,\gamma}.$
2.5. The geometric classification of compatible pre-Lie algebras
@@
$\overline{\mathcal{O}(\mathcal{C}_{28})},$  $\overline{\mathcal{O}(\mathcal{C}_{33})},$  $\overline{\mathcal{O}(\mathcal{C}_{37}^{\alpha})},$  $\overline{\mathcal{O}(\mathcal{C}_{09}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{11}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{22}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{27}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{36}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})},$  $\overline{\mathcal{O}(\mathcal{C}_{41}^{\alpha,\beta})},$ $\overline{\mathcal{O}(\mathcal{C}_{24}^{\alpha,\beta,\gamma})},$  $\overline{\mathcal{O}(\mathcal{C}_{31}^{\alpha,\beta,\gamma})},$  $\overline{\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})}$ and  $\overline{\mathcal{O}(\mathcal{C}_{40}^{\alpha,\beta,\gamma})}.$
@@
Thanks to Theorem 7, we have the algebraic classification of $2$-dimensional compatible pre-Lie algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}(\mathcal{C}_{24}^{\alpha,\beta,\gamma})=\dim\mathcal{O}(\mathcal{C}_{31}^{\alpha,\beta,\gamma})=\dim\mathcal{O}(\mathcal{C}_{38}^{\alpha,\beta,\gamma})=\dim\mathcal{O}(\mathcal{C}_{40}^{\alpha,\beta,\gamma})$ $=$ $7,$
$\dim\mathcal{O}(\mathcal{C}_{09}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{11}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{22}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{27}^{\alpha,\beta})$ $=$
$\dim\mathcal{O}(\mathcal{C}_{36}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{39}^{\alpha,\beta})=\dim\mathcal{O}(\mathcal{C}_{41}^{\alpha,\beta})$ $=$ $6,$
$\dim\mathcal{O}(\mathcal{C}_{37}^{\alpha})$ $=$ $5$
$\dim\mathcal{O}(\mathcal{C}_{28})=\dim\mathcal{O}(\mathcal{C}_{33})$ $=$ $4$
All necessary reasons for non-degenerations are listed below.
@@
$\begin{array}[]{llll}\mathcal{C}_{24}^{*}\end{array}$$\not\rightarrow$ $\begin{array}[]{llll}\mathcal{C}_{22}^{\alpha,\beta},\ \mathcal{C}_{26}^{\alpha},\ \mathcal{C}_{27}^{\alpha,\beta},\ \mathcal{C}_{28}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}c_{21}^{\prime 1}=c_{21}^{\prime 2}=c_{22}^{1}=c_{22}^{2}=c_{22}^{\prime 1}=c_{22}^{\prime 2}=0\par\end{array}\right\}$
$\begin{array}[]{llll}\mathcal{C}_{31}^{*}\end{array}$$\not\rightarrow$ $\begin{array}[]{llll}\mathcal{C}_{33},\ \mathcal{C}_{36}^{\alpha,\beta},\ \mathcal{C}_{37}^{\alpha}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}c_{22}^{\prime 1}=c_{23}^{\prime 2}=c_{22}^{1}=c_{22}^{2}=0\par\end{array}\right\}$
$\begin{array}[]{llll}\mathcal{C}_{36}^{*}\end{array}$$\not\rightarrow$ $\begin{array}[]{llll}\mathcal{C}_{37}^{\alpha}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}c_{22}^{1}=c_{12}^{1}=c_{12}^{\prime 1}=0\par\end{array}\right\}$

### v2

@@ 3.2 The geometric classification of compatible commutative
Theorem 3.7.
The variety of complex $2$-dimensional compatible commutative associative algebras has dimension $7$ and it has $2$ irreducible components defined by $\overline{\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}}$ and $\overline{\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}}$. In particular, there are no rigid algebras in this variety.
@@
Thanks to Theorem 2.8, we have the algebraic classification of $2$-dimensional compatible commutative associative algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have $\dim\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}=7$ and $\dim\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}=6$. $\mathcal{C}_{38}^{*}\not\to\mathcal{C}_{39}^{\alpha,\beta}$ due to the following relation:
=$\mathcal{R}=\big{\{}c_{22}^{1}=c_{12}^{1}=0,\,c_{22}^{2}c_{12}^{\prime 2}+c_{12}^{2}c_{21}^{\prime 1}+c_{11}^{1}c_{22}^{\prime 2}=c_{11}^{\prime 1}c_{22}^{2}+c_{11}^{1}c_{21}^{\prime 1}+c_{12}^{2}c_{22}^{\prime 2}\big{\}}.$
Thanks to [25], we have $\mathcal{C}_{07}\to\big{\{}\mathcal{C}_{03},\mathcal{C}_{05}^{0},\mathcal{C}_{06}^{1}\big{\}}.$ All necessary degenerations are given by
$\mathcal{C}_{38}^{0,0,t^{-1}}\xrightarrow{(te_{1},te_{2})}\mathcal{C}_{07},\quad\mathcal{C}_{39}^{-it^{-1},it^{-1}}\xrightarrow{(ite_{1}-ite_{2},-t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{13}^{1},$
$\mathcal{C}_{38}^{0,0,\alpha}\xrightarrow{(ite_{1}-ite_{2},-t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{15}^{\alpha},\quad\mathcal{C}_{38}^{\frac{i}{2t},0,\frac{i}{t}}\xrightarrow{(ite_{1}-ite_{2},-t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{16}^{0},$
$\mathcal{C}_{38}^{-\frac{1+\alpha t^{2}}{4t^{2}},-\frac{1+\alpha t^{2}}{4t^{2}},\frac{\alpha t^{2}-3}{4t^{2}}}\xrightarrow{(ite_{1}-ite_{2},-t^{2}e_{1}-t^{2}e_{2})}\mathcal{C}_{18}^{\alpha},\quad\mathcal{C}_{38}^{0,-t,\beta+t}\xrightarrow{(e_{1},te_{2})}\mathcal{C}_{24}^{0,\beta,0},$
$\mathcal{C}_{38}^{0,0,\beta}\xrightarrow{(e_{1},te_{2})}\mathcal{C}_{25}^{0,\beta,0},\quad\mathcal{C}_{39}^{\alpha,t^{-1}}\xrightarrow{(e_{1},te_{2})}\mathcal{C}_{29}^{\alpha},\quad\mathcal{C}_{39}^{-t+\beta,\beta}\xrightarrow{(e_{1}+e_{2},te_{2})}\mathcal{C}_{31}^{1,\beta,\beta},$
$\mathcal{C}_{38}^{-\alpha-\beta t+t^{-1},-\beta t,t^{-1}}\xrightarrow{(e_{1}+e_{2},te_{2})}\mathcal{C}_{30}^{\alpha,\beta},\quad\mathcal{C}_{38}^{0,\beta,0}\xrightarrow{(e_{1}+e_{2},te_{2})}\mathcal{C}_{32}^{1,\beta,\beta},$
$\mathcal{C}_{38}^{-t^{-2},\alpha-(1+\beta t)t^{-2},(1+\beta t)t^{-2}}\xrightarrow{(e_{1}+e_{2},te_{2})}\mathcal{C}_{34}^{\alpha,\beta},\quad\mathcal{C}_{38}^{\frac{1+2\alpha t}{2t},\alpha,0}\xrightarrow{(e_{1}+e_{2},-te_{1}+te_{2})}\mathcal{C}_{35}^{\alpha}.$
3.3 The geometric classification of compatible associative algebras
@@
Thanks to Theorem 2.10, we have the algebraic classification of $2$-dimensional compatible associative algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}=7,\quad\dim\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}=6,\quad\dim\mathcal{O}(\mathcal{C}_{28})=\dim\mathcal{O}(\mathcal{C}_{34})=4.$
Thanks to Theorem 3.8, we have that $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{39}^{\alpha,\beta}$ give irreducible components in the variety of $2$-dimensional compatible commutative associative algebras. The rest of the necessary degenerations are given by
$\mathcal{C}_{28}\xrightarrow{(e_{2},te_{1})}\mathcal{C}_{05}^{1},\quad\mathcal{C}_{33}\xrightarrow{(e_{2},te_{1})}\mathcal{C}_{06}^{0},\quad\mathcal{C}_{28}\xrightarrow{(e_{1}+\beta e_{2},te_{2})}\mathcal{C}_{25}^{1,\beta,\beta},\quad\mathcal{C}_{33}\xrightarrow{(e_{1}+\beta e_{2},te_{2})}\mathcal{C}_{32}^{0,\beta,0}.$
3.4 The geometric classification of compatible Novikov algebras
@@
Theorem 3.9.
The variety of complex $2$-dimensional compatible Novikov algebras has dimension $7$ and it has $6$ irreducible components defined by $\overline{\mathcal{O}(\mathcal{C}_{33})}$, $\overline{\mathcal{O}\big{(}\mathcal{C}_{09}^{\alpha,\beta}\big{)}}$, $\overline{\mathcal{O}\big{(}\mathcal{C}_{22}^{\alpha,\beta}\big{)}}$, $\overline{\mathcal{O}\big{(}\mathcal{C}_{39}^{\alpha,\beta}\big{)}}$, $\overline{\mathcal{O}\big{(}\mathcal{C}_{31}^{\alpha,\beta,\gamma}\big{)}}$ and $\overline{\mathcal{O}\big{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\big{)}}$. In particular, there is one rigid algebra in this variety.
@@
Thanks to Theorem 2.12, we have the algebraic classification of $2$-dimensional compatible Novikov algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}=7,\quad\dim\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{09}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{22}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{31}^{\alpha,\beta,\gamma}\bigr{)}=6,$
$\dim\mathcal{O}(\mathcal{C}_{33})=4.$
Thanks to Theorem 3.9, we have that $\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{39}^{\alpha,\beta}$ give irreducible components in the variety of $2$-dimensional compatible commutative associative algebras. Algebras $\mathcal{C}_{09}^{\alpha,\beta}$, $\mathcal{C}_{22}^{\alpha,\beta}$, and $\mathcal{C}_{31}^{\alpha,\beta,\gamma}$ have a one-dimensional subalgebra with zero multiplication (concerning both multiplications), but $\mathcal{C}_{33}$ has not. Hence, $\mathcal{C}_{33}$ is not in the orbit closure of $\mathcal{C}_{09}^{\alpha,\beta}$, $\mathcal{C}_{22}^{\alpha,\beta}$, and $\mathcal{C}_{31}.$ The rest of the necessary degenerations are given by
$\mathcal{C}_{31}^{t,\frac{t+2}{2t},-\frac{1}{2}}\xrightarrow{(te_{1}+\frac{2t}{3}e_{2},\frac{t^{2}}{2}e_{1}+t^{2}e_{2})}\mathcal{C}_{01},\quad\mathcal{C}_{31}^{0,0,t^{-1}}\xrightarrow{(t^{2}e_{1}+e_{2},te_{1}-t^{2}e_{2})}\mathcal{C}_{04},$
$\mathcal{C}_{31}^{0,t^{-1},\alpha t^{-1}}\xrightarrow{(te_{1}-\alpha^{-1}t^{2}e_{2},te_{2})}\mathcal{C}_{06}^{\alpha\neq 0},\quad\mathcal{C}_{09}^{\frac{(2\alpha+t)\beta}{2\beta+t},-\frac{\beta t}{2\beta+t}}\xrightarrow{(\frac{2\beta+t}{2\beta}e_{1}-\frac{(2\beta+t)^{2}}{2\beta t}e_{2},\frac{(2\beta+t)t}{4\beta^{2}}te_{1})}\mathcal{C}_{10}^{\alpha,\beta\neq 0},$
$\mathcal{C}_{31}^{1,\alpha t^{-1},t^{-1}}\xrightarrow{(te_{1}-t^{2}e_{2},-t^{3}e_{2})}\mathcal{C}_{13}^{\alpha},\quad\mathcal{C}_{31}^{1,t^{-1},0}\xrightarrow{(te_{1}+\alpha^{-1}te_{2},\alpha^{-1}t^{2}e_{2})}\mathcal{C}_{14}^{\alpha\neq 0},$
$\mathcal{C}_{31}^{t^{-1}-t,0,-1+\alpha t^{-1}}\xrightarrow{(t^{2}e_{1}+\frac{t^{2}}{1-\alpha t}e_{2},te_{1}+\frac{t^{3}}{1-\alpha t}e_{2})}\mathcal{C}_{20}^{\alpha,0},\quad\mathcal{C}_{31}^{t^{-1}-t,0,-\alpha t^{-1}}\xrightarrow{(t^{2}e_{1}-\alpha^{-1}te_{2},te_{1}-\alpha^{-1}te_{2})}\mathcal{C}_{21}^{\alpha,0},$
$\mathcal{C}_{31}^{t^{-1}-t,\alpha t^{-1},\beta t^{-1}}\xrightarrow{(t^{2}e_{1}-\beta^{-1}te_{2},te_{1}-\beta^{-1}t^{2}e_{2})}\mathcal{C}_{23}^{\alpha,\beta},\quad\mathcal{C}_{31}^{\alpha,\beta,\gamma}\xrightarrow{(e_{1},t^{-1}e_{2})}\mathcal{C}_{32}^{\alpha,\beta,\gamma}.$
3.5 The geometric classification of compatible pre-Lie algebras
@@
$\overline{\mathcal{O}(\mathcal{C}_{28})},\quad\overline{\mathcal{O}(\mathcal{C}_{33})},\quad\overline{\mathcal{O}(\mathcal{C}_{37}^{\alpha})},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{09}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{11}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{22}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{27}^{\alpha,\beta}\bigr{)}},$
$\overline{\mathcal{O}\bigl{(}\mathcal{C}_{36}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{41}^{\alpha,\beta}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{24}^{\alpha,\beta,\gamma}\bigr{)}},\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{31}^{\alpha,\beta,\gamma}\bigr{)}},$
$\overline{\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}}\quad\text{and}\quad\overline{\mathcal{O}\bigl{(}\mathcal{C}_{40}^{\alpha,\beta,\gamma}\bigr{)}}.$
@@
Thanks to Theorem 2.7, we have the algebraic classification of $2$-dimensional compatible pre-Lie algebras. After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}\bigl{(}\mathcal{C}_{24}^{\alpha,\beta,\gamma}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{31}^{\alpha,\beta,\gamma}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{38}^{\alpha,\beta,\gamma}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{40}^{\alpha,\beta,\gamma}\bigr{)}=7,$
$\dim\mathcal{O}\bigl{(}\mathcal{C}_{09}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{11}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{22}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{27}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{36}^{\alpha,\beta}\bigr{)}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{39}^{\alpha,\beta}\bigr{)}$
${}=\dim\mathcal{O}\bigl{(}\mathcal{C}_{41}^{\alpha,\beta}\bigr{)}=6,$
$\dim\mathcal{O}\bigl{(}\mathcal{C}_{37}^{\alpha}\bigr{)}=5,\quad\dim\mathcal{O}(\mathcal{C}_{28})=\dim\mathcal{O}(\mathcal{C}_{33})=4.$
All necessary reasons for non-degenerations are listed as follows:
@@
$\mathcal{C}_{24}^{*}\not\rightarrow\mathcal{C}_{22}^{\alpha,\beta},\,\mathcal{C}_{26}^{\alpha},\,\mathcal{C}_{27}^{\alpha,\beta},\,\mathcal{C}_{28}$ =$\mathcal{R}=\bigl{\{}c_{21}^{\prime 1}=c_{21}^{\prime 2}=c_{22}^{1}=c_{22}^{2}=c_{22}^{\prime 1}=c_{22}^{\prime 2}=0\bigr{\}}$
$\mathcal{C}_{31}^{*}\not\rightarrow\mathcal{C}_{33},\,\mathcal{C}_{36}^{\alpha,\beta},\,\mathcal{C}_{37}^{\alpha}$ =$\mathcal{R}=\bigl{\{}c_{22}^{\prime 1}=c_{23}^{\prime 2}=c_{22}^{1}=c_{22}^{2}=0\bigr{\}}$
$\mathcal{C}_{36}^{*}\not\rightarrow\mathcal{C}_{37}^{\alpha}$ =$\mathcal{R}=\bigl{\{}c_{22}^{1}=c_{12}^{1}=c_{12}^{\prime 1}=0\bigr{\}}$